# Non-Resolution Reasoning (NRR): A Computational Framework for Contextual Identity and Ambiguity Preservation

### v4

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We demonstrate NRR’s advantages through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide a minimal empirical validation on a synthetic context-shift task where an NRR-lite model achieves 90.9% out-of-distribution accuracy compared to 9.1% for standard architectures, demonstrating that ambiguity preservation enables structural generalization. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@
Keywords: Non-resolution reasoning, contextual identity, ambiguity preservation, multi-vector embeddings, non-collapsing attention, structural generalization
@@ 1.2 Contributions of This Work
    Empirical Demonstration: We provide a minimal empirical validation on a synthetic context-shift task where NRR-lite achieves 90.9% out-of-distribution accuracy compared to 9.1% for standard architectures.
@@ 5.4 Strategic Resolution: When and How to Collapse
6 Empirical Validation
To demonstrate the $A\neq A$ principle, we implemented a minimal experiment validating the claim: maintaining contextual non-identity until sufficient context arrives improves generalization.
6.1 Task Design: Context-Switch Word Sense (CSWS)
We created a synthetic two-turn task where the identity of an ambiguous token (“bank”) is determined only after context arrives.
  • Turn 1 (Ambiguous): “The bank is {adj}.” (e.g., solid, stable)
  • Turn 2 (Context): Disambiguating cue (e.g., “investor” or “ducks”).
Task: Classify the sense of “bank” (FINANCIAL vs RIVER). Turn 1 provides insufficient information, forcing models to either commit early (risk error) or preserve ambiguity.
@@ 6.2 Model Comparison
  • Baseline Transformer: Standard single embedding per token.
  • NRR-lite: Maintains $k=2$ separate embeddings for the ambiguous token. Uses context gating in Turn 2 to resolve the embedding.
Note: NRR-lite is not a full implementation of NRR; it isolates the ambiguity-preservation component to demonstrate the causal effect of delayed resolution on generalization.
@@ 6.3 Results
Table 1: Minimal Experiment Results: Context-Switch Word Sense (CSWS) Task
Model ID Acc OOD Acc $H(g)$ Turn 1 $H(g)$ Turn 2
Baseline 1.000 0.091
NRR-lite 0.920 0.909 0.693 0.692
Figure 2: Minimal Experiment Results. Left: Ambiguity preservation measured by gate entropy $H(g)$. NRR-lite maintains high entropy at Turn 1 (no context). Right: Generalization performance. NRR-lite achieves 90.9% OOD accuracy while Baseline fails at 9.1%, demonstrating that ambiguity preservation enables structural rather than surface-level learning.
@@
    Baseline Failure (9.1% OOD): The baseline memorized specific cue words but failed to generalize to unseen cues.
@@
    NRR Success (90.9% OOD): By maintaining ambiguity at Turn 1 and resolving based on Turn 2 context structure, NRR-lite learned a transferable representation.
@@
    Ambiguity Preservation: NRR-lite maintained maximum entropy ($\approx\log 2$) at Turn 1, confirming it did not commit prematurely.
This experiment validates that maintaining contextual non-identity ($A\neq A$) enables structural generalization where standard architectures fail.
@@ 7 Future Evaluation Framework
Comprehensive evaluation requires scaling beyond the minimal experiment. We propose benchmarks for:
@@ 8 Conclusion
For readers interested in the theoretical underpinnings and philosophical framework of the proposed Non-Resolution Reasoning mechanism, please refer to our companion paper [34]. This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. We demonstrated through theory and minimal experimentation that preserving ambiguity is not a cost but a capability. Our experiment showed that delaying resolution allows for 90.9% generalization accuracy in ambiguous contexts where standard models fail (9.1%). NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.

### v5

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We demonstrate NRR’s advantages through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.63$, near-maximum $0.69$) at the ambiguous turn, while standard architectures collapse early ($H=0.10$), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@
Keywords: Non-resolution reasoning, contextual identity, ambiguity preservation, multi-vector embeddings, non-collapsing attention, entropy measurement
@@ 1.2 Contributions of This Work
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.63$ (near-maximum) while baselines collapse to $H=0.10$, robust across multiple random seeds.
@@ 5.4 Strategic Resolution: When and How to Collapse
6 Functional Verification
Scope of this section: This section does not aim to establish empirical superiority of NRR over existing large language models. Instead, we provide a minimal functional verification (proof-of-concept) demonstrating that the proposed architecture behaves as intended: it preserves ambiguity before contextual cues arrive and resolves it afterward. Comparative evaluation against production LLMs is reserved for future work.
To demonstrate the $A\neq A$ principle, we designed a verification procedure that directly measures ambiguity preservation: does the model maintain interpretive uncertainty when context is insufficient?
6.1 Task Design: Turn 1 Entropy Measurement
We created a synthetic two-turn disambiguation task where the identity of an ambiguous token (“bank”) is determined only after context arrives.
  • Turn 1 (Ambiguous): “The bank is {adj}.” (e.g., solid, stable, old, new)
  • Turn 2 (Context): Disambiguating cue (e.g., “investor” $\to$ FINANCIAL, “ducks” $\to$ RIVER).
Key Insight: Rather than measuring classification accuracy (which depends on label distribution), we directly measure entropy of the output distribution at Turn 1—before any disambiguating context arrives.
$H(p)=-\sum_{i}p_{i}\log p_{i}$ (4)
For a binary classification:
  • $H_{\max}=\log 2\approx 0.693$ (perfect ambiguity: $p=[0.5,0.5]$)
  • $H_{\min}=0$ (complete certainty: $p=[1,0]$ or $[0,1]$)
A model that preserves ambiguity should maintain high entropy at Turn 1; a model that collapses early will show low entropy.
@@ 6.2 Model Comparison
  • Baseline: Standard single embedding per token, 32-dimensional embeddings, 2-layer MLP classifier.
  • NRR-lite: Maintains $k=2$ separate embeddings for the ambiguous token (“bank”). A context-dependent gate selects the appropriate embedding based on Turn 2 context. When Turn 2 is neutral (no context), the gate remains balanced.
Both models were trained on 1,000 samples with balanced labels (50% FINANCIAL, 50% RIVER) for 100 epochs.
@@ 6.3 Results
Table 1: Turn 1 Entropy Verification Results (5 random seeds)
Model Turn 1 Entropy $H$ Gate Entropy Context Accuracy
Baseline $0.102\pm 0.089$ 100%
NRR-lite $\mathbf{0.629\pm 0.025}$ $0.693$ 100%
Maximum $H$ $0.693$ $0.693$
Figure 2: Turn 1 Entropy Verification Results. Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.63$), near the theoretical maximum, while Baseline collapses ($H\approx 0.10$). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1. NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 0 or 1 (committed). Right: Both models achieve 100% accuracy when Turn 2 context is provided, confirming that NRR-lite’s ambiguity preservation does not impair resolution capability.
@@
    Baseline Collapse ($H=0.10$): The baseline commits to one interpretation at Turn 1, even without disambiguating context. This premature collapse is the architectural bias NRR addresses.
@@
    NRR-lite Preservation ($H=0.63$): NRR-lite maintains near-maximum entropy at Turn 1, confirming that multiple interpretations coexist until context arrives.
@@
    Gate Behavior: NRR-lite’s context gate achieves $H=0.693$ (perfect balance) when no context is provided, shifting appropriately when Turn 2 arrives.
  4. 4.
    Resolution Capability: Both models achieve 100% accuracy with context, demonstrating that ambiguity preservation does not impair eventual resolution.
  5. 5.
    Robustness: Results replicate across 5 random seeds ($t=13.3$, $p<0.001$).
This verification confirms that NRR-lite implements the $A\neq A$ principle: the same symbol (“bank”) maintains distinct potential identities until context determines which applies.
@@ 7 Future Evaluation Framework
Comprehensive evaluation requires scaling beyond this minimal verification. We propose benchmarks for:
@@ 8 Conclusion
Theoretical foundations and philosophical framework will be addressed in a forthcoming companion paper. This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. We demonstrated through theory and functional verification that preserving ambiguity is not a cost but a capability. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.63$) when context is absent, while standard architectures collapse prematurely (H=0.10H=0.10)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.